OpenStudy (anonymous):

How do I find the vertex of -4.9t^2+32t+2 Im having a hard time factoring out the 4.9

jimthompson5910 (jim_thompson5910):

use the formula x = -b/(2a) where in this case b = 32 and a = -4.9

jimthompson5910 (jim_thompson5910):

this will give you the x coordinate of the vertex

jimthompson5910 (jim_thompson5910):

or in this case, the t coordinate once you figure that out, you plug it into the expression to find the y value of the vertex

OpenStudy (anonymous):

ok make sense. How would I do it using difference of squares

jimthompson5910 (jim_thompson5910):

this isn't a difference of squares i think you mean "completing the square" right?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

well there are 2 basic ways

OpenStudy (anonymous):

Im stuck on the decimal

jimthompson5910 (jim_thompson5910):

first is using the formula above to find the x coordinate of the vertex, then using that to find the y coordinate of the vertex that will give you h and k respectively which you can use to plug into y = a(x-h)^2 + k

jimthompson5910 (jim_thompson5910):

or you can do it like this

OpenStudy (anonymous):

I am trying to get it into vertex form like posted above

jimthompson5910 (jim_thompson5910):

-4.9t^2+32t+2 -4.9(t^2+32/(-4.9)t+2/(-4.9)) -4.9(t^2-6.5306122t-0.408163265) -4.9(t^2-6.5306122t + (6.5306122/2)^2 - (6.5306122/2)^2 -0.408163265) -4.9( (t^2-6.5306122t + (6.5306122/2)^2 - (6.5306122/2)^2) -0.408163265) -4.9( (t-(6.5306122/2))^2 - (6.5306122/2)^2 -0.408163265) -4.9( (t-(6.5306122/2))^2 -10.25406066) -4.9(t-(6.5306122/2))^2 -4.9*(-10.25406066) -4.9(t-(6.5306122/2))^2 + 50.244897234 -4.9(t-3.2653061)^2 + 50.244897234

jimthompson5910 (jim_thompson5910):

that's the formal way to complete the square...which you can see is very ugly

OpenStudy (isaiah.feynman):

Don't use the difference method so you don't get in trouble with decimals use the first method.

jimthompson5910 (jim_thompson5910):

so I recommend you use the first method

jimthompson5910 (jim_thompson5910):

and I made a typo somewhere, not sure where but you should end up with -4.9*(t-3.265306122)^2+54.24489795

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

here's how I would do it

jimthompson5910 (jim_thompson5910):

x = -b/(2a) x = -32/(2*(-4.9)) x = 3.265306122 so the x coordinate of the vertex is 3.265306122

jimthompson5910 (jim_thompson5910):

plug this into the function to find the y coordinate of the vertex y = -4.9t^2+32t+2 y = -4.9(3.265306122)^2+32(3.265306122)+2 y = 54.24489795 so the vertex is the point (3.265306122, 54.24489795)

OpenStudy (anonymous):

awsome.Thanks

jimthompson5910 (jim_thompson5910):

which means (h,k) = (3.265306122, 54.24489795) ---> h = 3.265306122, k = 54.24489795

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (jhannybean):

\begin{align} -4.9t^2+32t+2 &= (-4.9t^2 +32t) +2 \\ &= -4.9\left(t^2 -\frac{320}{49}t -\frac{25600}{2401} \right) +2 \\ &= -4.9\left(t-\frac{160}{49}\right)^2 +2 +\frac{2560}{49} \\ &= -4.9\left(t-\frac{160}{49}\right)^2 +\frac{2658}{49} \end{align} Vertex form : \(y=a(x-h)^2 +k\) your form: \(\large y= -4.9\left(x-\frac{160}{49}\right)^2 +\frac{2658}{49}\)

OpenStudy (unklerhaukus):

The quadratic\[ax^2+bx+c=0\] has a vertex with x-coordinate at \[x=\frac{-b}{2a}\] ______ The t-coordinate of the vertex of \[-4.9t^2+32t+2=0\] is at \[t=\frac{-32}{2\times-4.9}\approx3.2\]